ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 248, 2022

Catanzaro, M. J., V. Y. Chernyak and J. R. Klein Fluctuations of cycles in a finite CW complex

315

Chen, B., Y. Shen and L. Zhao Kähler Finsler metrics and conformal deformations

355

Chernikov, A. and K. Gannon Definable convolution and idempotent Keisler measures

271

Chernyak, V. Y. See M. J. Catanzaro Crone, L., L. Fishman and S. Jackson Hausdorff dimension

regularity properties and games

481

Cúth, M., M. Doucha and O. Kurka Complexity of distances: Reductions of distances between metric and Banach spaces

383

Doucha, M. See M. Cúth Feigin, E., A. Khoroshkin and I. Makedonskyi Duality theorems

for current groups

441

Fishman, L. See L. Crone Fox, J., J. Tidor and Y. Zhao Induced arithmetic removal:

complexity 1 patterns over finite fields

1

Gannon, K. See A. Chernikov Jackson, S. See L. Crone Ji, K. See C. Jiang Jiang, C., K. Ji and J. Wu Similarity invariants of essentially

normal Cowen–Douglas operators and Chern polynomials

229

Khoroshkin, A. See E. Feigin Klein, J. R. See M. J. Catanzaro Kurka, O. See M. Cúth Kutsenko, I. See C. Leon Lakrec, T. Scenery reconstruction for random walk on random

scenery systems

149

Leon, C., I. Kutsenko and V. Volpert Existence of solutions for a nonlocal reaction-diffusion equation in biomedical applications

67

Macias Castillo, D. and K.-W. Tsoi On non-abelian higher special elements of p-adic representations

95

Makedonskyi, I. See E. Feigin Makhul, M., O. Roche-Newton, S. Stevens and A. Warren

The Elekes–Szabó problem and the Uniformity Conjecture

39

Monderer, T. and D. Neftin Symmetric Galois groups under specialization

201

Neftin, D. See T. Monderer Roche-Newton, O. See M. Makhul Shen, Y. See B. Chen Stevens, S. See M. Makhul Tidor, J. See J. Fox Tsoi, K.-W. See D. Macias Castillo Volpert, V. See C. Leon Warren, A. See M. Makhul Wu, J. See C. Jiang Zhao, L. See B. Chen Zhao, Y. See J. Fox